# Broadening Discovery through Structural Models: Multimodal Combination of Local and Structural Properties for Predicting Chemical Features

### v1

@@ 4.3 Graph model
Creation and augmentation of graph.
A graph is constructed from SMILES representations utilizing the RDKit package, wherein each atom is represented as a vertex. Two parameters – atom number and chirality – are designated as attributes of the vertices. In this framework, each bond is represented as an edge, with the bond multiplicity (single, double, triple, or aromatic) serving as the attribute for the edges.
Subsequently, 20% of the atomic attributes are masked, replacing them with a designated mask token. In the case of graphomers, an equivalent approach is applied where 20% of the edge attributes are also masked, transforming these attributes into the mask token.
The augmentation process and the graph model operation scheme are shown in Figure  3.
@@
Model training.
In the graph component of our model, we have experimented with three distinct architectures: Graph Isomorphism Network (GIN), Graph Convolutional Network (GCN), and Graphormer. We employ augmentation techniques to transform the molecular graph into two distinct representations. Following this, we train the GCN, GIN, or Graphormer models with the objective of minimizing the differences between the augmentations of one graph and maximize difference between augmentations of different graphs (this process for graphs in one batch is shown in Figure 3).
@@ 4.4 Connection Between Models
Let $e_{\text{graph}}$, denote the output of the graph model and $e_{\text{lang}}$ represent the output of the language model. Furthermore, let $\psi_{\text{graph}}$ and $\psi_{\text{lang}}$ be the respective projection blocks for the graph and language models. Define $\mathbb{A}$ as the latent space of the graph model, $\mathbb{B}$ as the latent space of the language model, and $\mathbb{C}$ as the space into which the embeddings are projected. Thus, we have $e_{\text{graph}}\in\mathbb{A}$, $e_{\text{lang}}\in\mathbb{B}$ with $\psi_{\text{graph}}:\mathbb{A}\rightarrow\mathbb{C}$ and $\psi_{\text{lang}}:\mathbb{B}\rightarrow\mathbb{C}$.
@@ 4.5 Loss functions
where $L_{\text{lang}}$ is the loss function of the language model, $L_{\text{graph}}$ is the loss function of the graph part of the model, and $L_{\text{bimodal}}$ is the embedding projection loss function from the graph and language models. Coefficients $\alpha$, $\beta$, and $\gamma$ are some constants which can be considered hyperparameters.
Language model loss.
$L_{\text{lang}}$ is calculated as regular Cross-Entropy applied to labels and predicted tokens of the language model.
Graph model loss.
$L_{\text{graph}}$ is defined as NTXent-Loss [Sohn, 2016] applied to the batch of augmented graphs’ embeddings and to the batch of original graphs’ embeddings. It tries to minimize the distance between augmented and original ones of the same index and distances others with different indices. NTXent-Loss calculates the cosine distance between two vectors and uses the temperature parameter to balance positive and negative pairs. Let $sim(u,v)$ denotes the cosine similarity between vectors $u$ and $v$. Then the loss function for a positive pair of examples (i, j) is as follows:
$(L_{\text{graph}})_{i,j}=-\log\left(\frac{e^{\text{sim}(u_{i},v_{j})/\tau}}{\sum_{k=1}^{N}e^{\text{sim}(u_{i},v_{k})/\tau}}\right),$ (2)
@@
Bimodal Loss.
The bimodal loss, denoted as $L_{\text{bimodal}}$, is defined also as the NTXent-Loss applied to the output embeddings generated by both the language model and the graph model within a given batch. This loss function aims to minimize the distance between the embeddings of the same index from both models while maximizing the distance between embeddings corresponding to different indices.
To achieve this, we employ two distinct projection blocks to convert the embeddings from the graph and language models into a unified third latent space. Utilizing a single projection block to transfer the embeddings from one model into the latent space of the other could inadvertently lead to the training of one model to mimic the behavior of the other. Such an outcome is undesirable, as the distinct functionalities of the models enhance the universal applicability of the bimodal architecture.
5 Experiments
@@ 5.1 Pretraining datasets and data preparation
We pretrain our model on parts of two different datasets: PubChem [Kim et al., 2023] and ZINC [Irwin et al., 2012]. Initially, the compounds in them are stored in SMILES format. By leveraging a more physics-based input format, namely ECFP, and employing one of the most sophisticated language models, we achieved a significant milestone: a language model (LM) trained from scratch on two subsets of the PubChem dataset, comprising 2.5 million and 10 million entries, respectively. This model exhibits performance comparable to those trained on the largest datasets within the field. The data preparation process could be divided into two main parts (as shown in Figure 5).
Figure 5: An example of overall molecular data preprocessing. ECFP and Graph representations are generated from the SMILES sequence and are feed-forwarded to the language model and graph model, respectively.
Language model data. We construct ECFP from the obtained SMILES of the molecule (the algorithm is given in Appendix LABEL:app), and then we mask $15\%$ of the elements in the obtained array (after performing the tokenization process and considering them as tokens).
Graph model data. We build a graph based on the SMILES of the molecule and then use its augmentation, which transforms it into two different molecule graphs. The augmentation process consists of masking $20\%$ of randomly chosen atom types (for GCN and GIN) and masking both $20\%$ of randomly chosen atom types and edges (for Graphormer). We mask only types of atoms and edges, not the edges and atoms themselves (as in the MolCLR approach [Wang et al., 2022]), due to the greater physicochemical validity of this method. For example, if we mask the red highlighted edge in Figure 6 in octyl formate (with the SMILES encoding CCCCCCCCOC(=O)), we get two existing compounds—heptane (CCCCCCC) and methyl formate (COC(=O)). Thus, the model learns to converge the embeddings of octyl formate and the total embedding of heptane and methyl formate, which is fundamentally incorrect.
Figure 6: Example of dropping edges problem.
@@ 5.2 QSAR tasks
For the zero-shot evaluation of our proposed architecture, we select a set of widely recognized cheminformatics benchmarks based on the quantitative structure-activity relationship (QSAR). Although specially designed descriptors often outperform transformer-based models in these contexts, the simplicity of these benchmarks allows us to assess the quality and versatility of our architecture without the confounding influence of large-scale superstructures typically encountered in more complex problem-solving scenarios.
We evaluate four distinct models: RoBERTa (denoted as ECFP-BERT), which is trained on ECFP; ECFP-BERT in conjunction with a Graph Isomorphism Network (GIN); ECFP-BERT combined with a Graph Convolutional Network (GCN); and ECFP-BERT integrated with Graphormer. The first three models utilize a dataset comprising 10 million entries sourced from the PubChem database, while the fourth model is trained on a smaller dataset of 1 million entries from the same source.
Table 1: Results for classification tasks. ROC-AUC metric (higher is better) for BBBP, Tox21, ClinTox, BACE, MUV and HIV datasets.
@@
BBBP Tox21 (NR-AR) ClinTox (FDA_APPROVED) ClinTox (CT_TOX) BACE MUV HIV
MolCLR(GCN) [Wang et al., 2022] 0.72 0.70 0.66 0.69 0.71 0.67 0.78
MolCLR(GIN) [Wang et al., 2022] 0.74 0.74 0.87 0.77 0.81 0.57 0.76
ChemBERTa [Chithrananda et al., 2020] 0.64 0.75 - 0.73 0.72 0.66 0.62
ECFP-BERT (ours) 0.82 0.71 0.71 0.69 0.73 0.61 0.65
BERT+GIN (ours) 0.88 0.79 0.88 0.71 0.79 0.70 0.74
BERT+GCN (ours) 0.85 0.79 0.71 0.69 0.73 0.64 0.73
BERT+Graphormer (ours) 0.77 0.74 0.87 0.78 0.75 0.71 0.81
Table 2: Results regression tasks, MAE (less is better) metric for QM7, QM8 and QM9 datasets. MSE for FreeSolv, ESOL and Lipo.

### v2

@@ 4.3 Graph model
Creation and augmentation of graph. A graph is constructed from SMILES representations utilizing the RDKit package, wherein each atom is represented as a vertex. Two parameters – atom number and chirality – are designated as attributes of the vertices. In this framework, each bond is represented as an edge, with the bond multiplicity (single, double, triple, or aromatic) serving as the attribute for the edges.
Subsequently, 20% of the atomic attributes are masked, replacing them with a designated mask token. In the case of graphomer, an equivalent approach is applied, where 20% of the edge attributes are also masked, transforming these attributes into the mask token. The augmentation process and the graph model operation scheme are shown in Figure  4.
@@
Model training. In the graph component of our model, we have experimented with three distinct architectures: Graph Isomorphism Network (GIN), Graph Convolutional Network (GCN), and Graphormer. We employ augmentation techniques to transform the molecular graph into two distinct representations. Following this, we train the GCN, GIN, or Graphormer models with the objective of minimizing the differences between the augmentations of one graph and maximizing the differences between augmentations of different graphs (this process for graphs in one batch is shown in Figure 4).
@@ 4.4 Connection Between Models
Let $e_{\text{graph}}$ denote the output of the graph model and $e_{\text{lang}}$ represent the output of the language model. Furthermore, let $\psi_{\text{graph}}$ and $\psi_{\text{lang}}$ be the respective projection blocks for the graph and language models. Define $\mathbb{A}$ as the latent space of the graph model, $\mathbb{B}$ as the latent space of the language model, and $\mathbb{C}$ as the space into which the embeddings are projected. Thus, we have $e_{\text{graph}}\in\mathbb{A}$, $e_{\text{lang}}\in\mathbb{B}$ with $\psi_{\text{graph}}:\mathbb{A}\rightarrow\mathbb{C}$ and $\psi_{\text{lang}}:\mathbb{B}\rightarrow\mathbb{C}$.
@@ 4.5 Loss functions
where $L_{\text{lang}}$ is the loss function of the language model, $L_{\text{graph}}$ is the loss function of the graph part of the model, and $L_{\text{bimodal}}$ is the embedding projection loss function from the graph and language models. The coefficients $\alpha$, $\beta$, and $\gamma$ are constants that can be considered hyperparameters and are assigned default values of $1.0$.
Language model loss. $L_{\text{lang}}$ is calculated as the regular Cross-Entropy applied to the labels and predicted tokens of the language model.
Graph model loss. $L_{\text{graph}}$ is defined as NTXent-Loss [Sohn, 2016] applied to the batch of augmented graphs’ embeddings and to the batch of original graphs’ embeddings. It tries to minimize the distance between augmented and original embeddings of the same index and distances others with different indices. NTXent-Loss calculates the cosine distance between two vectors and utilizes the temperature parameter to balance the influence of positive and negative pairs. Let $sim(u,v)$ denotes the cosine similarity between vectors $u$ and $v$. Then, the loss function for a positive pair of examples (i, j) is as follows:
$(L_{\text{graph}})_{i,j}=-\log\left(\tfrac{e^{\text{sim}(u_{i},v_{j})/\tau}}{\sum_{k=1}^{N}e^{\text{sim}(u_{i},v_{k})/\tau}}\right),$ (2)
@@
Bimodal Loss. The bimodal loss, denoted as $L_{\text{bimodal}}$, is defined also as the NTXent-Loss applied to the output embeddings generated by both the language model and the graph model within a given batch. This loss function aims to minimize the distance between the embeddings of the same index from both models while maximizing the distance between embeddings corresponding to different indices. To achieve this, we employ two distinct projection blocks to convert the embeddings from the graph and language models into a unified third latent space. Utilizing a single projection block to transfer the embeddings from one model into the latent space of the other could inadvertently lead to the training of one model to mimic the behavior of the other. Such an outcome is undesirable, as the distinct functionalities of the models enhance the universal applicability of the bimodal architecture.
5 Experiments and limitations
Figure 6: Example of dropping edges problem.
@@ 5.1 Pretraining datasets and data preparation
We pretrain our model on parts of the PubChem [Kim et al., 2023] dataset. Initially, the compounds in them are stored in SMILES format. By splitting into chemically relevant substructures, leveraging a more physics-based input format (descriptors) and employing one of the most sophisticated language models, we achieve a significant milestone: a language model (LM) trained from scratch with 10 million entries from the PubChem dataset. The pre-training process takes 98 hours for RoBERTa, about 250 hours for BERT+GCN and BERT+GIN, and approximately 400 hours for BERT+Graphormer (each result on the A100 GPU for 10 epochs). This model exhibits performance comparable to those trained on the largest datasets within the field. The data preparation process can be divided into two main parts.
Language model data. We construct the descriptors’ sequence as it was described in previous paragraphs, and then we mask $15\%$ of the elements in the obtained array (after performing the tokenization process and considering them as tokens).
Graph model data. We build a graph based on the SMILES of the molecule and then use its augmentation, which transforms it into two different molecule graphs. The augmentation process consists of masking $20\%$ randomly chosen atom types (for GCN and GIN) and masking both $20\%$ randomly chosen atom types and edges (for Graphormer). We mask only types of atoms and edges, not the edges and atoms themselves (as in the MolCLR approach [Wang et al., 2022]), due to the greater physicochemical validity of this method. For example, if we mask the red highlighted edge in Figure 6 in octyl formate (with the SMILES encoding CCCCCCCCOC(=O)), we obtain two existing compounds: heptane (CCCCCCC) and methyl formate (COC(=O)). Thus, the model learns to converge the embeddings of octyl formate and the total embedding of heptane and methyl formate, which is fundamentally incorrect.
@@ 5.2 QSAR tasks
For zero-shot evaluation, we select widely recognized cheminformatics benchmarks focused on quantitative structure-activity relationship (QSAR) tasks. Although specially designed descriptors often outperform transformer models in this context, the simplicity of these benchmarks allows for the assessment of our architecture’s quality and versatility without the influence of large-scale superstructures in complex problems.
We assess four models: RoBERTa (denoted SubD-BERT), trained on descriptors; SubD-BERT combined with a Graph Isomorphism Network (GIN); SubD-BERT paired with a Graph Convolutional Network (GCN); and SubD-BERT integrated with Graphormer. All models utilize a dataset of 10 million entries from PubChem.
Additionally, classical machine learning models (XGBoost, LightGBM, SVM) trained on the same descriptors provide strong baselines, known for QSAR effectiveness with handcrafted features. It can be seen that even such simple models show quite good results on many presented benchmarks, which demonstrates the correctness of the chosen paradigm of partitioning the molecule into substructures and further construction of descriptors.
Table 1: Results for classification tasks. ROC-AUC metric (higher is better) for BBBP, Tox21, ClinTox, BACE, MUV and HIV datasets, the scaffold split was used for train-test-validation (80-10-10) split.
@@
BBBP Tox21 (NR-AR) ClinTox (FDA APPROVED) ClinTox (CT TOX) BACE MUV HIV
MolCLR(GCN) [Wang et al., 2022] 0.723±0.025 0.704±0.002 0.668±0.035 0.694±0.032 0.711±0.090 0.676±0.019 0.787±0.005
MolCLR(GIN) [Wang et al., 2022] 0.742±0.020 0.740±0.003 0.872±0.031 0.775±0.037 0.814±0.07 0.796±0.017 0.761±0.006
ChemBERTa [Chithrananda et al., 2020] 0.647±0.053 0.753±0.009 - 0.736±0.015 0.721±0.022 0.667±0.015 0.625±0.012
Uni-Mol [Zhou et al., 2023] 0.729±0.006 0.796±0.005 0.895±0.018 0.711±0.023 0.857±0.002 0.821±0.013 0.808±0.003
GEM [Fang et al., 2022] 0.724±0.004 0.781±0.001 0.875±0.013 0.692±0.019 0.856±0.011 0.817±0.005 0.806±0.009
GROVER (base) [Rong et al., 2020] 0.700±0.001 0.743±0.001 0.812±0.030 0.664±0.032 0.826±0.007 0.673±0.018 0.625±0.009
GROVER (large) [Rong et al., 2020] 0.695±0.001 0.735±0.001 0.75±0.037 0.683±0.041 0.810±0.014 0.673±0.018 0.682±0.011
Molformer [Wu et al., 2023] 0.916±0.002 - 0.907±0.006 0.812±0.031 0.844±0.017 - -
MolFormer-XL [Ross et al., 2022] 0.917±0.001 0.847±0.001 0.933±0.004 0.901±0.012 0.862±0.009 - 0.812±0.003
SubD-BERT (ours) 0.893±0.018 0.829±0.007 0.947±0.013 0.926±0.017 0.811±0.022 0.753±0.015 0.692±0.011
BERT+GIN (ours) 0.937±0.002 0.852±0.003 0.912±0.009 0.924±0.014 0.855±0.015 0.832±0.011 0.786±0.007
BERT+GCN (ours) 0.891±0.005 0.830±0.002 0.903±0.016 0.793±0.031 0.738±0.012 0.794±0.017 0.736±0.010
BERT+Graphormer (ours) 0.862±0.009 0.815±0.003 0.878±0.019 0.837±0.021 0.892±0.015 0.819±0.013 0.851±0.060
XGBoost (descriptors, ours) 0.821 0.663 0.856 0.871 0.695 0.650 0.562
LightGBM (descriptors, ours) 0.832 0.653 0.886 0.853 0.682 0.581 0.546
SVM (descriptors, ours) 0.612 0.617 0.525 0.679 0.547 0.559 0.534
Table 2: Results regression tasks, MAE (less is better) metric for QM7, QM8 and QM9 datasets. MSE for FreeSolv, ESOL and Lipo, the scaffold split was used for train-test-validation (80-10-10) split.